# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ VII $l$-Dependence of Correlations Between Order-Parameter Measures
Figure 23: Plots of the correlation coefficient $r$ between $\Delta_{T}$ and $\Delta_{S}$ (orange), $\Delta_{S}$ and $\Delta_{L}$ (blue), and $\Delta_{T}$ and $\Delta_{L}$ (purple) as a function of patch size $l$ for (left to right) various values of the average order parameter. The dashed lines indicate the values of $r$ for the ”$l=0$” case where the bond order parameters are independently and identically randomly distributed. The same $x$-axis scale is used in all three plots.
@@
Figure 24: Plots of the correlation coefficient $r$ between $\Delta_{S}$ and $\rho_{C}$ (blue), and between $\Delta_{T}$ and $\rho_{C}$ (purple) as a function of patch size $l$ for (left to right) various values of the average order parameter. The dashed lines indicate the values of $r$ for the ”$l=0$” case where the bond order parameters are independently and identically randomly distributed. The same $x$-axis scale is used in all three plots.
@@
In this section, we fully quantify the observations made in the preceding sections about the correlations present among various order-parameter measures ($\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$) and the coherence-peak height ($\rho_{C}$) and examine these as a function of the inhomogeneity length scale $l$. To assess the correlation between a pair of variables, we use the Pearson correlation coefficient $r$: a perfectly linear correlation between two variables gives $r=1$, the absence of any correlation gives $r=0$, and a perfectly linear anticorrelation results in $r=-1$. As such, this quantity is an excellent measure of the accuracy of the order-parameter measurables ($\Delta_{S}$ and $\Delta_{L}$) as compared to the true underlying order parameter ($\Delta_{T}$). To obtain $r$, we collect all values of the relevant pair of quantities measured on all sites and across four disorder realizations, giving us $1,048,576$ data points per quantity.
@@
In Fig. 24, we plot $r$ as a function of $l$ for the following pairs of quantities: $\Delta_{S}$-$\Delta_{T}$, $\Delta_{S}$-$\Delta_{L}$, and $\Delta_{T}$-$\Delta_{L}$. We consider once more three values of the mean order parameter $\overline{\Delta}$: $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$. For the $\Delta_{S}$-$\Delta_{T}$ pair, the behavior of $r$ is to a good extent independent of $\overline{\Delta}$. The correlation can be seen to be very small when $l\ll\xi$. At the smallest value of $l/\xi$ considered ($l=1$ for $\overline{\Delta}=0.08$, corresponding to $l\approx 0.11\xi$), $r$ for this pair of variables is only around 0.3. (This correlation is actually even smaller for the maximally disordered “$l=0$” case where the bond order parameters are independently distributed, for which the correlation drops to around 0.2.) When $l<\xi$, the coefficient generally is less than 0.65, but in the physically interesting regime where $l\approx\xi$, $r$ falls within the range $r\in[0.6,0.9]$, with the correlation being weaker for $\overline{\Delta}=0.08$ and stronger for $\overline{\Delta}=0.32$. When $l\gtrapprox 4\xi$, the correlation between this pair of variables is very strong, with $r>0.9$. That the correspondence between $\Delta_{S}$ and $\Delta_{T}$ improves with increasing $l/\xi$ is not surprising, but it is reassuring that in the regime relevant to the cuprates ($l\approx\xi$), $\Delta_{S}$ is a fairly accurate, if imperfect, proxy for the true order parameter when order-parameter inhomogeneity is present.
@@
Meanwhile, the correspondence between $\Delta_{L}$ and $\Delta_{T}$ is considerably weaker compared to that between the $\Delta_{S}$-$\Delta_{T}$ pair, except when $l/\xi<2$ for $\overline{\Delta}=0.08$ and $l/\xi<0.5$ for $\overline{\Delta}=0.16$. While $r$ increases monotonically with $l$ in all cases, we find that $r$ at large values of $l/\xi$ (such as $l/\xi\approx 8$) is still below the value of $0.9$ seen in the $\Delta_{S}$-$\Delta_{T}$ case (at which point $r$ will have saturated to almost unity in that case). The difference between $r$ for the $\Delta_{L}$-$\Delta_{T}$ pair and that for the $\Delta_{S}$-$\Delta_{T}$ pair is most pronounced in the $\overline{\Delta}=0.32$ case. We find additionally that $r$ generally depends on $l$, but not on $l/\xi$: regardless of the value of $\overline{\Delta}$ used, the obtained $r$ remains largely the same for fixed $l$. When $l=1$, $r\approx 0.6$, and this value barely changes until $l=8$ is reached, at which point $r$ increases. But even for $l=32$, $r\approx 0.8$, suggesting an imperfect correspondence even at very large patch sizes. This is presumably due to the lingering presence of kinks even at large patch sizes which continue to homogenize the slope of the low-energy density of states, making $\Delta_{L}$ more uniform than the underlying order parameter $\Delta_{T}$ even within this regime. Similar statements can be made about the correlation of the $\Delta_{S}$-$\Delta_{L}$ pair: $r$ in this case is consistently smaller than that of the $\Delta_{S}$-$\Delta_{T}$ case (except for a narrow window of $l$ values in the $\overline{\Delta}=0.08$ case). Also, $r$ appears to depend on $l$ alone and not on $l/\xi$. Interestingly, $r$ for $\Delta_{S}$-$\Delta_{L}$ and $r$ for $\Delta_{T}$-$\Delta_{L}$ differ from one another at small values of $l$, but start to match almost exactly when $l\geq 8$. The reason for this is the fact that both $\Delta_{S}$ and $\Delta_{L}$ are very narrowly distributed for small $l$ and their joint distribution consequently appears as a small, highly localized ellipse (see Fig. 6c for an illustration), leading to a small $r$ in this regime.
@@
The correlation coefficients for $\Delta_{S}$-$\rho_{C}$ and $\Delta_{T}$-$\rho_{C}$ pairs behave similarly as a function of $l/\xi$ for all values of $\overline{\Delta}$ considered. When $l/\xi<4$, $r$ for both pairs is negative, and only when $l/\xi$ is increased beyond this threshold does the correlation become positive. In the $l\approx\xi$ regime, $r$ for the $\Delta_{S}$-$\rho_{C}$ pair generally lies between $-0.3$ and $-0.6$, while $r$ for $\Delta_{T}$-$\rho_{C}$ falls within the $[-0.4,-0.7]$ range. Strong correlations ($r\approx 0.7$ and above) between $\rho_{C}$ and either order-parameter measure only start to appear at very large patch sizes with $l/\xi\gtrapprox 32$; the fact that the correspondence remains imperfect even with large patches suggests that the bulk regime is still not fully realized, and that boundary effects between patches play a role in spoling the expected strong correlation between the coherence-peak height and the order parameter in the large-patch regime.
@@ VIII.1 The Width of the Distribution of the Order-Parameter Inhomogeneity
These results open up the possibility that the cuprates are even more inhomogeneous than presently supposed from estimates taken from the width of the spectral-gap distribution. Across a broad doping range (from $p=0.08$ to $p=0.22$), it is found that the standard deviation $\sigma$ of $\Delta_{S}$ for Bi-2212 is $\sigma\approx 0.2\overline{\Delta_{S}}$ McElroy et al. (2005a); Alldredge et al. (2008); Gomes et al. (2007); Pasupathy et al. (2008). If we suppose that the underlying order-parameter ($\Delta_{T}$) inhomogeneity in the cuprates is in the $l\approx\xi$ regime, then $\sigma$ is an underestimate of the width of the $\Delta_{T}$ distribution. In particular, it is conceivable that there exist patches where the pairing interaction vanishes but the local spectral gap is nonzero due to the proximity effect.
@@
No direct comparison of the distributions of $\Delta_{S}$ and $\Delta_{T}$ has been performed thus far for the cuprates, although it is now to obtain a direct image of the underlying order parameter via Josephson STS methods, which use superconducting flakes as tips in the scanning tunneling microscope Hamidian et al. (2016); Cho et al. (2019); Du et al. (2020). A local measurement of the critical current $I_{c}$ using these Josephson tunnel-junction techniques turns out to be a faithful representation of $\Delta_{T}$ Graham and Morr (2017, 2019); Sulangi et al. (2021). That said, such a comparison has already been made in an iron-based superconductor (FeTe0.55Se0.55), where it was found that the $I_{c}$ distribution (and consequently that of $\Delta_{T}$) is considerably broader than that of the spectral gap Cho et al. (2019). While there are differences between that system and the cuprates (e.g., the pairing symmetry of FeSeTe is different from that of the cuprates, and stronger chemical-potential disorder is present in FeSeTe), this particular material provides a concrete example of a system for which the width of the spectral-gap distribution severely underestimates that of the order-parameter distribution.
@@
In the absence of a comparison between experimentally obtained $\Delta_{S}$ and $\Delta_{T}$, the correlations we obtained involving the coherence-peak height can be used as a guide to estimate the length scale of the underlying order-parameter inhomogeneity. STS experiments generally find that $\Delta_{S}$ and $\rho_{C}$ are negatively correlated with one another across all dopings where superconductivity is present. In our simulations, we find this to be true when $l<4\xi$, regardless of the value of $\overline{\Delta}$ used; this provides an upper bound for the length scale of the $\Delta_{T}$ inhomgeneity. This is further evidence for the $l\approx\xi$ scenario that is believed to be the case in the cuprates across a broad doping range.
@@
We finally note that in the $l\approx\xi$ regime, the real-space map of $\Delta_{S}$ (Fig. 8b) looks very similar to the various gap maps from various STS experiments in the cuprates (e.g., Fig. 1 of Fang et al. Fang et al. (2006), Fig. 1 of McElroy et al. McElroy et al. (2005a), and Fig. 5a of Ye et al. Ye et al. (2024)), the square shape of the underlying patches notwithstanding. This demonstrates that the underlying order parameter can, at least in principle, look very different from the spectral gap map. In our model, we assumed that the patches in $\Delta_{T}$ are all of the same size, yet the resulting $\Delta_{S}$ map shows that the regions with the smallest gaps have the largest spatial extent and vice versa, similar to what is seen in STS (e.g., Fang et al. Fang et al. (2006)). Our results suggest the possibility that the small- and large-$\Delta_{T}$ regions all have the same typical size: the $\Delta_{S}$ measurements provide an overestimate and an underestimate of the spatial extent of the small- and large-$\Delta_{T}$ patches, respectively.
@@ VIII.2 Origin of Kinks and Low-Energy Homogeneity
Figure 25: Top left: Plots of the binned average differential conductance $dI/dV$ taken from an underdoped ($T_{c}=79$ K, $p\approx 15$%) Bi-2212 sample. Reprinted figure with permission from McElroy et al., Phys. Rev. Lett 94 197005 (2005) McElroy et al. (2005a). Copyright (2005) by the American Physical Society. Top right, bottom left, and bottom right: Plots of the binned average local density of states taken from simulated data with $\Delta=0.32$, $l=4$ (b); $\Delta=0.16$, $l=8$ (c); and $\Delta=0.08$, $l=16$ (d), respectively. All three of these scenarios fall within the $l\approx\xi$ regime.
@@
Consistently seen in STS studies of many cuprates are kinks in the differential conductance spectra, which appear at a characteristic energy $\Delta_{K}$ that is parametrically less than $\Delta_{S}$. In underdoped-to-optimally doped Bi-2212, these kinks are ubiquitous and prominent, and found to be present mainly in large-gap regions McElroy et al. (2005a); Alldredge et al. (2008); Pushp et al. (2009); Alldredge et al. (2013). For overdoped Bi-2212, these are also present in the spectra, albeit in more muted form, with the kink energy closer to the average spectral gap Pushp et al. (2009); Alldredge et al. (2013). Meanwhile, in Bi-2201, these kinks remain prominent seen even at moderate overdoping, close to the end of the superconducting dome Ye et al. (2024). The origin of these kinks is not fully known, and it is generally believed that since these kinks are more prominent at small dopings, these are an effect of strong correlations. However, our work provides a clear picture of their origin from a mean-field perspective: these arise from proximity coupling between patches with different order-parameter values. This proximity effect is strongest at low energies, where the quasiparticle response is very similar to that of a homogeneous $d$-wave superconductor with order parameter $\overline{\Delta}$. This homogeneity is lost past the kink energy $\Delta_{K}$, which serves as a dividing line between the homogeneous states and the inhomogeneous ones. The kinks therefore arise from the system finally starting to reflect the underlying inhomogeneity for $\omega>\Delta_{K}$ and deviating from the homogeneous lineshapes at low-energies. This also explains why the coherence-peak heights are inversely correlated with the spectral gap: much of the spectral weight that would have been in the large-gap coherence peak in the absence of inhomogeneity is put instead in the kink, resulting in a shorter coherence peak. Note that this mechanism does not invoke a frequency-independent self-energy that arises separately from inhomogeneity (e.g., from inelastic scattering). We will return to this point later.
@@
This mesoscopic superconducting proximity-effect-based mechanism appears to explain well the kinks seen in both Bi-2212 and Bi-2201 across a broad doping range from underdoping to overdoping. In particular, Fig. 9a resembles very closely Fig. 2a of McElroy et al. McElroy et al. (2005a), which describes bin-averaged spectra taken from a mildly underdoped ($T_{c}=79$ K) Bi-2212 sample. We provide a side-by-side comparison of the McElroy et al. results with a more extensive selection of our simulated data (with $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$) in Fig. 25. Fig. 25d also looks very similar to Figs. 1e and 1f of Ye et al. Ye et al. (2024), which describes binned LDOS spectra for optimally doped and mildly overdoped Bi-2201. (One can observe from Fig. 25d that the kinks are less prominent the smaller $\overline{\Delta}$ is, a likely explanation for their decreasing prominence in the overdoped regime.) For these two particular experimental datasets, the kink energies appear to coincide with the smallest spectral gap, which supports the proximity-effect mechanism outlined earlier.
@@
However, at strong underdoping, the mechanism outlined here is arguably too simplistic to account for the particularly severe kinks seen in this regime. In particular, Pushp et al. Pushp et al. (2009) and McElroy et al. McElroy et al. (2005a) both find that nearly all spectra in this regime show kinks. Within our model, some of the spectra (the small-gap ones) should not exhibit kinks at all, so this effect at strong underdoping appears to lie beyond our mechanism. This likely signals a breakdown of the mean-field $d$-wave description in this regime. We hypothesize that the kinks at severe underdoping arise not only from the superconducting proximity effect, but also from strong correlation effects resulting from proximity to the nearby antiferromagnetic insulating phase, which together have the effect of “pushing out” the spectral gaps as frequency is increased, resulting in most spectra having kinks as opposed to only the large-gap ones.

### v2

@@ VII $l$-Dependence of Correlations Between Order-Parameter Measures
Figure 22: Plots of the correlation coefficient $r$ between $\Delta_{T}$ and $\Delta_{S}$ (orange), $\Delta_{S}$ and $\Delta_{L}$ (blue), and $\Delta_{T}$ and $\Delta_{L}$ (purple) as a function of patch size $l$ for various values of the average order parameter. The dashed lines indicate the values of $r$ for the ”l=0l=0” case where the bond order parameters are independently and identically randomly distributed. The same $x$-axis scale is used in all three plots.
@@
Figure 23: Plots of the correlation coefficient $r$ between $\Delta_{S}$ and $\rho_{C}$ (blue), and between $\Delta_{T}$ and $\rho_{C}$ (purple) as a function of patch size $l$ for various values of the average order parameter. The dashed lines indicate the values of $r$ for the ”l=0l=0” case where the bond order parameters are independently and identically randomly distributed. The same $x$-axis scale is used in all three plots.
@@
In this section, we fully quantify the observations made in the preceding sections about the correlations present among various order-parameter measures ($\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$) and the coherence-peak height ($\rho_{C}$) and examine these as a function of the inhomogeneity length scale $l$. To assess the correlation between a pair of variables, we use the Pearson correlation coefficient $r$: a perfectly linear correlation between two variables gives $r=1$, the absence of any correlation gives $r=0$, and a perfectly linear anticorrelation results in r=1r=-1. As such, this quantity is an excellent measure of the accuracy of the order-parameter measurables ($\Delta_{S}$ and $\Delta_{L}$) as compared with the true underlying order parameter ($\Delta_{T}$). To obtain $r$, we collect all values of the relevant pair of quantities measured on all sites and across four disorder realizations, giving us $1,048,576$ data points per quantity.
@@
In Fig. 22, we plot $r$ as a function of $l$ for the following pairs of quantities: $\Delta_{S}$-$\Delta_{T}$, $\Delta_{S}$-$\Delta_{L}$, and $\Delta_{T}$-$\Delta_{L}$. We consider once more three values of the mean order parameter $\overline{\Delta}$: $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$. For the $\Delta_{S}$-$\Delta_{T}$ pair, the behavior of $r$ is to a good extent independent of $\overline{\Delta}$. The correlation can be seen to be very small when $l\ll\xi$. At the smallest value of $l/\xi$ considered ($l=1$ for $\overline{\Delta}=0.08$, corresponding to $l\approx 0.11\xi$), $r$ for this pair of variables is only around 0.3. (This correlation is actually even smaller for the maximally disordered “l=0l=0” case where the bond order parameters are independently distributed, for which the correlation drops to around 0.2.) When $l<\xi$, the coefficient generally is less than 0.65, but in the physically interesting regime where $l\approx\xi$, $r$ falls within the range $r\in[0.6,0.9]$, with the correlation being weaker for $\overline{\Delta}=0.08$ and stronger for $\overline{\Delta}=0.32$. When $l\gtrapprox 4\xi$, the correlation between this pair of variables is very strong, with $r>0.9$. That the correspondence between $\Delta_{S}$ and $\Delta_{T}$ improves with increasing $l/\xi$ is not surprising, but it is reassuring that in the regime relevant to the cuprates ($l\approx\xi$), $\Delta_{S}$ is a fairly accurate, if imperfect, proxy for the true order parameter when order-parameter inhomogeneity is present.
@@
Meanwhile, the correspondence between $\Delta_{L}$ and $\Delta_{T}$ is considerably weaker compared with that between the $\Delta_{S}$-$\Delta_{T}$ pair, except when $l/\xi<2$ for $\overline{\Delta}=0.08$ and $l/\xi<0.5$ for $\overline{\Delta}=0.16$. While $r$ increases monotonically with $l$ in all cases, we find that $r$ at large values of $l/\xi$ (such as $l/\xi\approx 8$) is still below the value of $0.9$ seen in the $\Delta_{S}$-$\Delta_{T}$ case (at which point $r$ will have saturated to almost unity in that case). The difference between $r$ for the $\Delta_{L}$-$\Delta_{T}$ pair and that for the $\Delta_{S}$-$\Delta_{T}$ pair is most pronounced in the $\overline{\Delta}=0.32$ case. We find additionally that $r$ generally depends on $l$, but not on $l/\xi$: regardless of the value of $\overline{\Delta}$ used, the obtained $r$ remains largely the same for fixed $l$. When $l=1$, $r\approx 0.6$, and this value barely changes until $l=8$ is reached, at which point $r$ increases. But even for $l=32$, $r\approx 0.8$, suggesting an imperfect correspondence even at very large patch sizes. This is presumably due to the lingering presence of kinks even at large patch sizes which continue to homogenize the slope of the low-energy density of states, making $\Delta_{L}$ more uniform than the underlying order parameter $\Delta_{T}$ even within this regime. Similar statements can be made about the correlation of the $\Delta_{S}$-$\Delta_{L}$ pair: $r$ in this case is consistently smaller than that of the $\Delta_{S}$-$\Delta_{T}$ case (except for a narrow window of $l$ values in the $\overline{\Delta}=0.08$ case). Also, $r$ appears to depend on $l$ alone and not on $l/\xi$. Interestingly, $r$ for $\Delta_{S}$-$\Delta_{L}$ and $r$ for $\Delta_{T}$-$\Delta_{L}$ differ from one another at small values of $l$, but start to match almost exactly when $l\geq 8$. The reason for this is the fact that both $\Delta_{S}$ and $\Delta_{L}$ are very narrowly distributed for small $l$ and their joint distribution consequently appears as a small, highly localized ellipse (see Fig. 5c for an illustration), leading to a small $r$ in this regime.
@@
The correlation coefficients for $\Delta_{S}$-$\rho_{C}$ and $\Delta_{T}$-$\rho_{C}$ pairs behave similarly as a function of $l/\xi$ for all values of $\overline{\Delta}$ considered (see Fig. 23). When $l/\xi<4$, $r$ for both pairs is negative, and only when $l/\xi$ is increased beyond this threshold does the correlation become positive. In the $l\approx\xi$ regime, $r$ for the $\Delta_{S}$-$\rho_{C}$ pair generally lies between 0.3-0.3 and 0.6-0.6, while $r$ for $\Delta_{T}$-$\rho_{C}$ falls within the [0.4,0.7][-0.4,-0.7] range. Strong correlations ($r\approx 0.7$ and above) between $\rho_{C}$ and either order-parameter measure only start to appear at very large patch sizes with $l/\xi\gtrapprox 32$; the fact that the correspondence remains imperfect even with large patches suggests that the bulk regime is still not fully realized, and that boundary effects between patches play a role in spoiling the expected strong correlation between the coherence-peak height and the order parameter in the large-patch regime.
@@ VIII.1 The Width of the Distribution of the Order-Parameter Inhomogeneity
These results open up the possibility that the cuprates are even more inhomogeneous than presently supposed from estimates taken from the width of the spectral-gap distribution. Across a broad doping range (from $p=0.08$ to $p=0.22$), it is found that the standard deviation $\sigma$ of $\Delta_{S}$ for Bi-2212 is $\sigma\approx 0.2\overline{\Delta_{S}}$ [6, 10, 9, 11]. If we suppose that the underlying order-parameter ($\Delta_{T}$) inhomogeneity in the cuprates is in the $l\approx\xi$ regime, then $\sigma$ is an underestimate of the width of the $\Delta_{T}$ distribution. In particular, it is conceivable that there exist patches where the pairing interaction vanishes but the local spectral gap is nonzero due to the proximity effect.
@@
No direct comparison of the distributions of $\Delta_{S}$ and $\Delta_{T}$ has been performed thus far for the cuprates, although it is now to obtain a direct image of the underlying order parameter via Josephson STS methods, which use superconducting flakes as tips in the scanning tunneling microscope [20, 66, 22]. A local measurement of the critical current $I_{c}$ using these Josephson tunnel-junction techniques turns out to be a faithful representation of $\Delta_{T}$ [67, 68, 44]. That said, such a comparison has already been made in an iron-based superconductor (FeTe0.55Se0.55), where it was found that the $I_{c}$ distribution (and consequently that of $\Delta_{T}$) is considerably broader than that of the spectral gap [66]. While there are differences between that system and the cuprates (e.g., the pairing symmetry of FeSeTe is different from that of the cuprates, and stronger chemical-potential disorder is present in FeSeTe), this particular material provides a concrete example of a system for which the width of the spectral-gap distribution severely underestimates that of the order-parameter distribution.
@@
In the absence of a comparison between experimentally obtained $\Delta_{S}$ and $\Delta_{T}$, the correlations we obtained involving the coherence-peak height can be used as a guide to estimate the length scale of the underlying order-parameter inhomogeneity. STS experiments generally find that $\Delta_{S}$ and $\rho_{C}$ are negatively correlated with one another across all dopings where superconductivity is present. In our simulations, we find this to be true when $l<4\xi$, regardless of the value of $\overline{\Delta}$ used; this provides an upper bound for the length scale of the $\Delta_{T}$ inhomogeneity. This is further evidence for the $l\approx\xi$ scenario that is believed to be the case in the cuprates across a broad doping range.
@@
We finally note that in the $l\approx\xi$ regime, the real-space map of $\Delta_{S}$ (Fig. 7b) looks very similar to the various gap maps from various STS experiments in the cuprates (e.g., Fig. 1 of Fang et al. [8], Fig. 1 of McElroy et al. [6], and Fig. 5a of Ye et al. [32]), the square shape of the underlying patches notwithstanding. This demonstrates that the underlying order parameter can, at least in principle, look very different from the spectral gap map. In our model, we assumed that the patches in $\Delta_{T}$ are all of the same size, yet the resulting $\Delta_{S}$ map shows that the regions with the smallest gaps have the largest spatial extent and vice versa, similar to what is seen in STS (e.g., Fang et al. [8]). Our results suggest the possibility that the small- and large-$\Delta_{T}$ regions all have the same typical size: the $\Delta_{S}$ measurements provide an overestimate and an underestimate of the spatial extent of the small- and large-$\Delta_{T}$ patches, respectively.
@@ VIII.2 Origin of Kinks and Low-Energy Homogeneity
Figure 24: Top left: Plots of the binned average differential conductance $dI/dV$ taken from an underdoped ($T_{c}=79$ K, $p\approx 15$%) Bi-2212 sample. Reprinted figure with permission from McElroy et al., Phys. Rev. Lett 94 197005 (2005) [6]. Copyright (2005) by the American Physical Society. Top right, bottom left, and bottom right: Plots of the binned average local density of states taken from simulated data with $\Delta=0.32$, $l=4$ (b); $\Delta=0.16$, $l=8$ (c); and $\Delta=0.08$, $l=16$ (d), respectively. All three of these scenarios fall within the $l\approx\xi$ regime.
@@
Consistently seen in STS studies of many cuprates are kinks in the differential conductance spectra, which appear at a characteristic energy $\Delta_{K}$ that is parametrically less than $\Delta_{S}$. In underdoped-to-optimally doped Bi-2212, these kinks are ubiquitous and prominent, and found to be present mainly in large-gap regions [6, 10, 12, 14]. For overdoped Bi-2212, these are also present in the spectra, albeit in more muted form, with the kink energy closer to the average spectral gap [12, 14]. Meanwhile, in Bi-2201, these kinks remain prominent seen even at moderate overdoping, close to the end of the superconducting dome [32]. The origin of these kinks is not fully known, and it is generally believed that since these kinks are more prominent at small dopings, these are an effect of strong correlations. However, our work provides a clear picture of their origin from a mean-field perspective: these arise from proximity coupling between patches with different order-parameter values. This proximity effect is strongest at low energies, where the quasiparticle response is very similar to that of a homogeneous $d$-wave superconductor with order parameter $\overline{\Delta}$. This homogeneity is lost past the kink energy $\Delta_{K}$, which serves as a dividing line between the homogeneous states and the inhomogeneous ones. The kinks therefore arise from the system finally starting to reflect the underlying inhomogeneity for $\omega>\Delta_{K}$ and deviating from the homogeneous lineshapes at low-energies. This also explains why the coherence-peak heights are inversely correlated with the spectral gap: much of the spectral weight that would have been in the large-gap coherence peak in the absence of inhomogeneity is put instead in the kink, resulting in a shorter coherence peak. Note that this mechanism does not invoke a frequency-independent self-energy that arises separately from inhomogeneity (e.g., from inelastic scattering). We return to this point later.
@@
This mesoscopic superconducting proximity-effect-based mechanism appears to explain well the kinks seen in both Bi-2212 and Bi-2201 across a broad doping range from underdoping to overdoping. In particular, Fig. 8a resembles very closely Fig. 2a of McElroy et al. [6], which describes bin-averaged spectra taken from a mildly underdoped ($T_{c}=79$ K) Bi-2212 sample. We provide a side-by-side comparison of the McElroy et al. results with a more extensive selection of our simulated data (with $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$) in Fig. 24. Fig. 24d also looks very similar to Figs. 1e and 1f of Ye et al. [32], which describes binned LDOS spectra for optimally doped and mildly overdoped Bi-2201. (One can observe from Fig. 24d that the kinks are less prominent the smaller $\overline{\Delta}$ is, a likely explanation for their decreasing prominence in the overdoped regime.) For these two particular experimental datasets, the kink energies appear to coincide with the smallest spectral gap, which supports the proximity-effect mechanism outlined earlier.
@@
However, at strong underdoping, the mechanism outlined here is arguably too simplistic to account for the particularly severe kinks seen in this regime. In particular, Pushp et al. [12] and McElroy et al. [6] both find that nearly all spectra in this regime show kinks. Within our model, some of the spectra (the small-gap ones) should not exhibit kinks at all, so this effect at strong underdoping appears to lie beyond our mechanism. This likely signals a breakdown of the mean-field $d$-wave description in this regime. We hypothesize that the kinks at severe underdoping arise not only from the superconducting proximity effect, but also from strong correlation effects resulting from proximity to the nearby antiferromagnetic insulating phase, which together have the effect of “pushing out” the spectral gaps as frequency is increased, resulting in most spectra having kinks as opposed to only the large-gap ones.